OpenStudy (anonymous):

Solve x2 − 8x + 5 = 0 using the completing-the-square method. x = negative four plus or minus the square root of eleven x = four plus or minus the square root of eleven x = negative four plus or minus the square root of five x = four plus or minus the square root of five

OpenStudy (anonymous):

The Answer is b right?

OpenStudy (anonymous):

dd

OpenStudy (misssunshinexxoxo):

Correct! :)

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

x = -b + - square root b^2 - 4ac/ 2a x = -8 + - square root 8^2 - 4 (1) (-5) / 2 (1) x = -8 + - square root 64 + 20/ 2 x = -8 + - square root 84/ 2 = -8 + - square root 4 * 21/ 2 = -8 + - 2 * square root 21 / 2 x = 2 ( -4 + - square root 21 ) / 2 x = -4 + - square root 21

OpenStudy (anonymous):

Using the completing-the-square method, find the vertex of the function f(x) = 5x2 + 10x + 8 and indicate whether it is a minimum or a maximum and at what point. Maximum at (1, 8) Minimum at (1, 8) Maximum at (–1, 3) Minimum at (–1, 3)

OpenStudy (anonymous):

I place D

OpenStudy (misssunshinexxoxo):

Yes correct again! :) I'm taking algebra 2 FLVS also

OpenStudy (anonymous):

haha what module

OpenStudy (misssunshinexxoxo):

In module 8 right now did it under 16 weeks

OpenStudy (anonymous):

when did you start? I started 2 weeks ago haha and I am on Module 3

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

Started last month feel free to message me

OpenStudy (anonymous):

Well I plan on finishing it during the summer What is the hardest module in your perspective?

OpenStudy (anonymous):

@misssunshinexxoxo